OpenStudy (anonymous):

What is the simplified form of the quantity 3 x minus 6 all over x minus 2 ?

OpenStudy (anonymous):

x-2 with a restriction of positive 3 is what i got

OpenStudy (kropot72):

Factorising the nemerator gives 3(x - 2). 3(x - 2)/(x -2)=3

OpenStudy (anonymous):

ok i see that but im not understanding the restriction part

OpenStudy (kropot72):

Where did you get the restriction requirement from?

OpenStudy (anonymous):

the question asks for it

OpenStudy (kropot72):

The answer to the question you asked is 3. Please give the full question.

OpenStudy (anonymous):

Example: What is the simplified form of the quantity 2 x minus 8 all over x minus 4? (2x - 8)/(x -4) you can common factor out the 2 in the numerator: (2x - 8) = 2(x -4) the new fraction: 2(x-4) / (x-4) (x - 4)s cancel: 2/1 =2